OpenStudy (anonymous):

how do i know arccos -1/sqrt2 = 3pi/4

OpenStudy (anonymous):

\[\theta=\cos^{-1} -1/\sqrt{2} \]

OpenStudy (anonymous):

there is a unit circle you should memorize. i'll attach a pic in a sec

OpenStudy (anonymous):

this is all stuff they will expect you to have memorized. well, the first and second quadrants atleast

OpenStudy (anonymous):

since arccos is the inverse of cosine, you can rewrite the equation as: cos(3pi/4) = -1/sqrt(2). And this is plainly true if you can quickly sketch the cosine function, which you should know.

OpenStudy (anonymous):

hm, maybe i should learn it then.... would you just do that by wrote learning? isn't there a slightly easier way (takes slightly less time than wrote?)

OpenStudy (anonymous):

the ordered pairs are (cos(theta),sin(theta))

OpenStudy (anonymous):

Yes, you should memorize it, though there is a very nice reason why these functions look the way they do.. I would recommend the khan-academy.org website for an in depth explanation

OpenStudy (anonymous):

i feel like i should know what an ordered pair is...

OpenStudy (anonymous):

ordered pair: (x,y)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

gah, khan's voice annoys me.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Paul Zord, what course are you taking?

OpenStudy (anonymous):

You never learned unit circle?

OpenStudy (anonymous):

its a first year subject

OpenStudy (anonymous):

i've learnt it and forgotten it.

OpenStudy (anonymous):

If you've learned it before, you can learn it again. It's easy....

OpenStudy (anonymous):

paul:one of the great things aboput open study is that it allows you to revise all your basic concepts from ...a time long long ago

OpenStudy (anonymous):

Unit circle is based on the fact that the radius of a circle is one and then you can find all of the other angles using triangles, special angles rules and trigonometry

OpenStudy (anonymous):

And coverting between radians and degrees is easy because pi = 180 for example: 3pi/4 = 3(180)/4 = 3(45) = 135 degrees

OpenStudy (anonymous):

radians = 135*(pi/180)